OpenStudy (k8lyn911):

Calculus II "Integrals of Trig Products" homework question **Due by midnight If the integral tan^2(x) sec(x) dx from 0 to pi/4 = I, express the value of tan^4(x) sec(x) dx in terms of I.

OpenStudy (xapproachesinfinity):

hint \[\tan^2x=\sec^2x-1\]

OpenStudy (k8lyn911):

I've got it down to \[\int\limits_{0}^{\pi/4} \tan^4 (x) \sec (x) dx = \tan^3(x)\sec(x) - 3\int\limits_{0}^{\pi/4} \sec^2 (x) \tan(x)dx\] using \[\int\limits_{}^{} udv = uv - \int\limits_{}^{} vdu\] but I don't know where to go from there.

OpenStudy (xapproachesinfinity):

hmm let see try \[\tan^4x\sec x=(\sec^2x-1)^2\sec x\]

OpenStudy (xapproachesinfinity):

give it a try and if you need further help tag me ^_-

OpenStudy (xapproachesinfinity):

it would work with IBP too but you did something wrong along the way

OpenStudy (xapproachesinfinity):

just be persistent to find the solution hehe don't give up easily :)

OpenStudy (k8lyn911):

I'm not sure that I can figure it out within the time limit.

OpenStudy (xapproachesinfinity):

\[\int \tan^4x\sec xdx=\int (\sec^2x-1)\tan^2x\sec xdx\] \[=\int \tan ^2x \sec^3 xdx-I\]

OpenStudy (xapproachesinfinity):

take it from there

OpenStudy (xapproachesinfinity):

i gotta leave to sleep

OpenStudy (k8lyn911):

Oh, okay. Thanks for the help. :) I'll work on it.

OpenStudy (rizags):

if you really need to have this answered, i believe i have it

OpenStudy (rizags):

But its been a while since calc II so i dont really know if its correct

OpenStudy (k8lyn911):

I'm out of time now, so figuring out how to get the answer is more important now.

OpenStudy (rizags):

ok, well here is my answer for a similar problem i did in calc ii

OpenStudy (rizags):

If I= integral from 0 to pi/4 of tan^6(x)sec(x)dx, then express the value of the integral from 0 to pi/4 of tan^8(x)sec(x)dx in terms of I.

OpenStudy (rizags):

my answer for this (very similar) problem, was as follows: ∫ [0,π/4] tan⁸ x sec x dx = ∫ [0,π/4] tan⁶ x tan² x sec x dx = ∫ [0,π/4] tan⁶ x ( sec² x - 1 ) sec x dx = ∫ [0,π/4] tan⁶ x sec³ x dx - ∫ [0,π/4] tan⁶ x sec x dx = ∫ [0,π/4] tan⁶ x sec³ x dx - I ............ from (1) ................... Ans.

OpenStudy (rizags):

this is NOT your problem, but it is similar

OpenStudy (k8lyn911):

The answer is supposed to be \[\frac{ \sqrt{2} }{ 4 } - \frac{ 3 }{ 4 } I\] but I haven't figured out how to get there.

OpenStudy (irishboy123):

no frills solution

OpenStudy (k8lyn911):

Thanks! :)